Given that z is a standard normal random variable, compute the following probabilities (to 4 decimals).

  1. P(-1.91 ≤ z ≤ 0.41)

  2. P(0.55 ≤ z ≤ 1.26)

  3. P(-1.75 ≤ z ≤ -1.08)
Answer #1

Answer :

(a). P(-1.91 ≤ z ≤ 0.41)

= p( z ≤ 0.41 ) - p( z ≤ -0.91 )

=> using the standard normal z table

= 0.6591 - 0.1814

= 0.4777

  P(-1.91 ≤ z ≤ 0.41) = 0.4777

(b). P(0.55 ≤ z ≤ 1.26)

= p( z ≤ 1.26 ) - p( z ≤ 0.55 )

=> using the standard normal z table

= 0.8961 - 0.7088

= 0.1873

  P(0.55 ≤ z ≤ 1.26) = 0.1873

(c). P(-1.75 ≤ z ≤ -1.08)

= p( z ≤ -1.08 ) - p( z ≤ -1.75 )

=>  using the standard normal z table

= 0.1400 - 0.0400

= 0.1000

  P(-1.75 ≤ z ≤ -1.08) = 0.1000